OpenStudy (buggiethebug):

what power is required to lift a 5.0 kg mass vertically with a constant speed of 3.0m/s

OpenStudy (anonymous):

i think, you can always easily solve units to get a formula P=E/t P=Watt=J/s J=N.m N=kg.m/s2 J=kg.m2/s2 P=kg.m2/s3 =kg.m/s2 x m/s = hence you get formula for P you can figure out the formula from here, now what you need to know is what F you should apply to carryout a vertical motion.

OpenStudy (anonymous):

\[P=Fv\] P=Power F=Force v=velocity

OpenStudy (anonymous):

in this case, the force is the weight of the object which you can find by multiplying the mass and the acceleration due to gravity (which is 9.81 approx.. 10) \[F=mg\]

OpenStudy (anoop27):

P=5*10*3=150 Ans